A pole-vaulter just clears the bar at 5.80 m and falls back to the ground. The change in the vaulter's potential energy during the fall is -3.70 x 10^3 J. what is his weight?
Answer #1
change in potential energy / the change in height.

3700 / 5.8 = 637.931034
=638 N
Answer #2
mg*(-5.8) = -3.7*1000
=> mg = 3700/(5.8) = 637.93N
Hence weight = 637.93N
Answer #3
638N